the graph of a function is on the interval [-3,3] and f(x)= x^2 +ax+b the absolute minimum is (-2,2) find f(1)
1
Expert's answer
2011-11-03T09:10:51-0400
-b/2c=-2
c=1 then -b=-4 then
b=4
4+2a+b=2
2a=-6
a=-3
x^2-3x+4
f(1)=1-3+4=2